# The Rational Torsion Subgroup of $J_{0}(\mathfrak{p}^{r})$

### v1

@@ 5 Two degeneracy maps
Fix a prime-power ideal $\mathfrak{n}=\mathfrak{p}^{r}$ of $A$ with $r\geq 2$. Recall that the curve $Y_{0}(\mathfrak{n})$ is the generic fibre of the coarse moduli scheme for the functor which associates to an $A$-scheme $S$ the set of isomorphism classes of pairs $(\phi,C_{\mathfrak{n}})$, where $\phi$ is a rank-$2$ Drinfeld module over $S$ with an $\mathfrak{n}$-cyclic subgroup $C_{\mathfrak{n}}$; for details, see [1], [3], and [7]. Let $\alpha_{\mathfrak{p}}(\mathfrak{n}),\beta_{\mathfrak{p}}(\mathfrak{n}):Y_{0}(\mathfrak{n})\rightarrow Y_{0}(\mathfrak{n}/\mathfrak{p})$ be the degeneracy maps with moduli-theoretic interpretations:
@@
where $C_{\mathfrak{n}/\mathfrak{p}}$ and $C_{\mathfrak{p}}$ are the subgroups of $C_{\mathfrak{n}}$ of order $\mathfrak{n}/\mathfrak{p}$ and $\mathfrak{p}$, respectively. These morphisms are proper, and hence uniquely extend to morphisms $\alpha,\beta:X_{0}(\mathfrak{n})\rightarrow X_{0}(\mathfrak{n}/\mathfrak{p})$; cf. [12, p. 584]. Moreover, by Picard functoriality, the maps $\alpha$ and $\beta$ induce two homomorphisms:
@@
Note that $U_{\mathfrak{p}}:=\alpha_{\mathfrak{p}}(\mathfrak{n})^{\ast}\circ\beta_{\mathfrak{p}}(\mathfrak{n})_{\ast}=\beta_{\mathfrak{p}}(\mathfrak{n}\mathfrak{p})_{\ast}\circ\alpha_{\mathfrak{p}}(\mathfrak{n}\mathfrak{p})^{\ast}$; cf. [17, Rem. 3.5]. The definition of $U_{\mathfrak{p}}$ here is the dual version of that in [12, p. 584], which are conjugate to each other by the Atkin-Lehner involution $W_{\mathfrak{n}}$; cf. [13, p. 398].
@@
For a prime ideal $\mathfrak{q}\lhd A$ not dividing $\mathfrak{n}$, the Hecke operator $T_{\mathfrak{q}}$ commutes with the Atkin–Lehner involution $W_{\mathfrak{n}}$, so $T_{\mathfrak{q}}$ coincides with its dual; cf. [12, p. 584] and [13, p. 398].
@@
Then for a cusp $\Gamma_{0}(\mathfrak{n})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i}\end{bmatrix}$ of $X_{0}(\mathfrak{n})_{\mathbb{C}_{\infty}}$ in the form defined in Sec. 2, we have
$\alpha(\Gamma_{0}(\mathfrak{n})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i}\end{bmatrix})=\begin{cases}\Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i}\end{bmatrix},&\text{if $0\leq i\leq r-1$.}\\ \\ \Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot\begin{bmatrix}1\\ \mathfrak{p}^{r-1}\end{bmatrix},&\text{if $i=r$.}\end{cases}$
$\beta(\Gamma_{0}(\mathfrak{n})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i}\end{bmatrix})=\begin{cases}\Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot\begin{bmatrix}0\\ 1\end{bmatrix},&\text{if $i=0$.}\\ \\ \Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i-1}\end{bmatrix},&\text{if $1\leq i\leq r$.}\end{cases}$
Moreover, the degeneracy coverings $\alpha$ and $\beta$ have degree $|\mathfrak{p}|$. From the above and [8, Lem. 2.2], $\alpha$ and $\beta$ also induce maps between the divisor groups:
@@
Recall that $(P(\mathfrak{n})_{d})$ is the sum of all the cusps of $X_{0}(\mathfrak{n})$ of height $d$. We have:
@@
Inspired from Ex. Example, we decompose
@@
The first statement of lemma follows directly from [8, Lem. 2.1]. By Lem. 4.4, we have $\beta_{\ast}(\widetilde{r}(\Delta_{\mathfrak{p}})):=\widetilde{r}(\Delta_{\mathfrak{p}})|U_{\mathfrak{p}}=|\mathfrak{p}|\cdot\widetilde{r}(\Delta)$. Hence, we derive that $\beta_{\ast}$ maps $\operatorname{div}(\Delta_{\mathfrak{p}})\in\operatorname{Div}(X_{0}(\mathfrak{n}))$ to $|\mathfrak{p}|\cdot\operatorname{div}(\Delta)\in\operatorname{Div}(X_{0}(\mathfrak{n}/\mathfrak{p}))$. From [8, Sec. 2.1], the divisor of $\Delta_{\mathfrak{p}}$ on $X_{0}(\mathfrak{n})$ is
@@
where $b_{j}\in\mathbb{N}$. Now, the result for $2\leq j\leq r$ follows from $\beta_{\ast}(P(\mathfrak{n})_{\mathfrak{p}^{j}})=\frac{b_{j-1}}{a_{j}}\cdot(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{j-1}})$. For $j=0\text{ or }1$, we have $\beta_{\ast}(P(\mathfrak{n})_{1})=x\cdot(P(\mathfrak{n}/\mathfrak{p})_{1})$ and $\beta_{\ast}(P(\mathfrak{n})_{\mathfrak{p}})=y\cdot(P(\mathfrak{n}/\mathfrak{p})_{1})$ for some $x,y\in\mathbb{N}$. Consider
$f:=\left(\frac{\Delta}{\Delta_{\mathfrak{p}}}\right)^{|\mathfrak{p}|}\left(\frac{\Delta_{\mathfrak{p}^{2}}}{\Delta_{\mathfrak{p}}}\right)\in\mathcal{E}_{\mathfrak{n}}.$
Then $\widetilde{r}(f)\in\mathcal{H}(\mathfrak{n},\mathbb{Z})$. According to Lem. 4.3 and 4.4, we have $\widetilde{r}(f)|U_{\mathfrak{p}}=0$. Moreover, referencing [8, Sec. 2.1], the divisor of $f$ on $X_{0}(\mathfrak{n})$ is given by:
@@
This implies that $y=\frac{|\mathfrak{p}|-1}{q-1}\cdot x$. Now, considering $a_{0}$, $a_{1}$, and $b_{0}$, we find
@@
Lemma 5.3 (cf. [17, Lem. 3.4]).
Fix a prime-power ideal $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ with $r\geq 2$. Let $\beta_{\ast}:\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{p}^{r}))(K)\rightarrow\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{p}^{r-1}))(K)$ be the map induced from Lem. 5.2. Then
@@
Let $C_{i}:=(P(\mathfrak{n})_{\mathfrak{p}^{i}})-\deg(P(\mathfrak{n})_{\mathfrak{p}^{i}})\cdot[\infty]\in\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$ for $10\leq i\leq r-$ 1, and $C_{i}^{\prime}:=(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{i}})-\deg(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{i}})\cdot[\infty]\in\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}/\mathfrak{p}))(K)$. Then $\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$ is generated by $\{C_{0},C_{1},\cdots,C_{r-1}\}$. By Lem. 5.2, we have
@@
6 Proof of Main Theorem
Fix a prime-power ideal $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ with $r\geq 2$. We prove Thm. 1.7 by induction on $r$. It is known that $\mathcal{C}(\mathfrak{p})=\mathcal{T}(\mathfrak{p})$. Assume that $\mathcal{C}(\mathfrak{p}^{r-1})_{\ell}=\mathcal{T}(\mathfrak{p}^{r-1})_{\ell}$ for a prime $\ell\nmid q(q-1)$. We want to show that $\mathcal{C}(\mathfrak{p}^{r})_{\ell}=\mathcal{T}(\mathfrak{p}^{r})_{\ell}$. Consider the following diagram:
@@ 6 Proof of Main Theorem
We call $U_{\mathfrak{p}}(\mathcal{T}(\mathfrak{p}^{r}))_{\ell}\subset\mathcal{T}(\mathfrak{p}^{r})_{\ell}$ the old part of $\mathcal{T}(\mathfrak{p}^{r})_{\ell}$. By Lem. 5.3 and induction hypothesis, we have
@@
From the above and [8, Lem. 3.6], the order of $\overline{C^{\prime}}$ in $\mathcal{C}(\mathfrak{p}^{r})$ divides $M(\mathfrak{p})|\mathfrak{p}|^{r-2}$. Since $\frac{\Delta_{\mathfrak{p}^{2}}}{\Delta_{\mathfrak{p}}}$ does not have a $p$-th root in $\mathcal{O}(\Omega)^{\ast}$ by [5, Cor. 3.18], we see that $|\mathfrak{p}|^{r-2}$ divides $\operatorname{ord}(\overline{C^{\prime}})$; cf. [8, Lem. 3.3]. Let $e:=\begin{pmatrix}\pi_{\infty}^{2}&\pi_{\infty}\\ 0&1\end{pmatrix}\in E(\mathscr{T})$; cf. [8, Sec. 2.2]. By [8, Lem. 2.4], we compute that
@@
which implies that $M(\mathfrak{p})$ also divides $\operatorname{ord}(\overline{C^{\prime}})$; cf. [8, Lem. 3.3]. Then we conclude that $\operatorname{ord}(\overline{C^{\prime}})=M(\mathfrak{p})|\mathfrak{p}|^{r-2}$.
@@
Combining the above with Thm. 1.4 and Lem. 5.3, the prime-to-$p$ part of $\mathcal{C}(\mathfrak{p}^{r})[U_{\mathfrak{p}}]$ is cyclic and generated by $|\mathfrak{p}|^{r-2}\cdot\overline{C^{\prime}}$ of order $M(\mathfrak{p})$. In contrast, the $p$-part of $\mathcal{C}(\mathfrak{p}^{r})[U_{\mathfrak{p}}]$ may not be cyclic.
@@
for a prime $\ell\nmid q(q-1)$ and large enough $k$. It remains to show that
$\mathcal{E}_{00}(\mathfrak{p}^{r},\mathbb{Z}/\ell^{k}\mathbb{Z})[U_{\mathfrak{p}}]\hookrightarrow\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}.$
@@
Recall that the order of $\overline{C^{\prime}}$ in $\mathcal{C}(\mathfrak{p}^{r})$ is $M(\mathfrak{p})|\mathfrak{p}|^{r-2}$. Then by Thm. 3.4 and Eq. (5), we have $E_{\mathfrak{n}}=M(\mathfrak{p})|\mathfrak{p}|^{r-2}\cdot\widetilde{r}\circ g(C^{\prime})\in\mathcal{H}(\mathfrak{n},\mathbb{Z})$.
@@
By Lem. 4.3 and 4.4, for any prime $\mathfrak{q}\lhd A$ not equal to $\mathfrak{p}$, we have $E_{\mathfrak{n}}|T_{\mathfrak{q}}=(|\mathfrak{q}|+1)E_{\mathfrak{n}}$ and $E_{\mathfrak{n}}|U_{\mathfrak{p}}=0$. Hence, $E_{\mathfrak{n}}$ is annihilated by both the actions of the Eisenstein ideal $\mathfrak{E}(\mathfrak{n})$ and the Hecke operator $U_{\mathfrak{p}}$. ∎
@@
This follows from Prop. 4.2 and the fact that $(f|T_{\mathfrak{m}})^{\ast}(1)=|\mathfrak{m}|f^{\ast}(\mathfrak{m})$; cf. [11, Lem. 3.2]. ∎
We provide an Atkin-Lehner type result similar to [14, Thm. 2.18]:
@@
By [14, Lem. 2.13], $h$ is $\Gamma_{0}(\mathfrak{n}/\mathfrak{p},\mathfrak{p})$-invariant and harmonic. Moreover, by Thm. 3.3, for any $e=\begin{pmatrix}\pi_{\infty}^{k}&y\\ 0&1\end{pmatrix}\in E(\mathscr{T})$ and $a\in A$, we have
@@
Note that $R\cdot E_{\mathfrak{n}}\subseteq\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$. It suffices to prove the reverse inclusion. By Eq. (2) and Lem. 3.6, we have
@@
First, by Lem. 6.4 and Thm. 6.5, since $\tilde{f}^{\ast}(1)=0$, there exists $h\in\mathcal{H}(\mathfrak{n}/\mathfrak{p},R)$ which is $\Gamma_{\infty}$-invariant such that $h|B_{\mathfrak{p}}=\tilde{f}$. Note that
@@
Second, by [12, Lem. 2.23], since $h$ is $\Gamma_{\infty}$-invariant, we have
@@
Note that
$\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\cdot[\infty]:=\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\cdot\begin{bmatrix}1\\ \mathfrak{n}\end{bmatrix}=\begin{bmatrix}1\\ \mathfrak{p}\end{bmatrix}$
on the cusps of $X_{0}(\mathfrak{n})$. Let $e_{j}:=\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix}\in E(\mathscr{T})$ with $j\geq 0$. Then $\{\Gamma_{0}(\mathfrak{n})\cdot e_{j}\mid j\geq 0\}$ forms an end on the quotient graph of the Bruhat-Tits tree $\Gamma_{0}(\mathfrak{n})\backslash\mathscr{T}$ corresponding to the cusp $\begin{bmatrix}1\\ \mathfrak{p}\end{bmatrix}$ on $X_{0}(\mathfrak{n})$; cf. [14, Def. 2.5] and [5, p. 184]. Before computing $\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]$, we consider the following lemma:
Lemma 6.7 (cf. [8, Lem. 2.4]).
@@
    $\widetilde{r}(\Delta)(\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix})=-(q-1)(q^{j+1}-q-1).$
@@
    $\widetilde{r}(\Delta_{\mathfrak{p}^{k}})(\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix})=-(q-1)(q^{j+1}|\mathfrak{p}|^{2-k}-q-1).$
@@
$\widetilde{r}(\Delta)(\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix})=\widetilde{r}(\Delta)(\begin{pmatrix}\pi_{\infty}^{-j}&0\\ \mathfrak{p}\pi_{\infty}^{-j}&1\end{pmatrix})$
@@
where the penultimate equality comes from [5, Eq. (2.3)].

### v2

@@ 5 Two degeneracy maps
Fix a prime power ideal $\mathfrak{n}=\mathfrak{p}^{r}$ of $A$ with $r\geq 2$. Recall that the curve $Y_{0}(\mathfrak{n})$ is the generic fibre of the coarse moduli scheme for the functor which associates to an $A$-scheme $S$ the set of isomorphism classes of pairs $(\phi,C_{\mathfrak{n}})$, where $\phi$ is a rank-$2$ Drinfeld module over $S$ with an $\mathfrak{n}$-cyclic subgroup $C_{\mathfrak{n}}$; for details, see [1], [3], and [8]. Let $\alpha_{\mathfrak{p}}(\mathfrak{n}),\beta_{\mathfrak{p}}(\mathfrak{n}):Y_{0}(\mathfrak{n})\rightarrow Y_{0}(\mathfrak{n}/\mathfrak{p})$ be the degeneracy maps with moduli-theoretic interpretations:
@@
where $C_{\mathfrak{n}/\mathfrak{p}}$ and $C_{\mathfrak{p}}$ are the subgroups of $C_{\mathfrak{n}}$ of order $\mathfrak{n}/\mathfrak{p}$ and $\mathfrak{p}$, respectively. These morphisms are proper, and hence uniquely extend to morphisms $\alpha,\beta:X_{0}(\mathfrak{n})\rightarrow X_{0}(\mathfrak{n}/\mathfrak{p})$; cf. [13, p. 584]. Moreover, by Picard functoriality, the maps $\alpha$ and $\beta$ induce two homomorphisms:
@@
Note that $U_{\mathfrak{p}}:=\alpha_{\mathfrak{p}}(\mathfrak{n})^{\ast}\circ\beta_{\mathfrak{p}}(\mathfrak{n})_{\ast}=\beta_{\mathfrak{p}}(\mathfrak{n}\mathfrak{p})_{\ast}\circ\alpha_{\mathfrak{p}}(\mathfrak{n}\mathfrak{p})^{\ast}$; cf. [18, Remark 3.5]. The definition of $U_{\mathfrak{p}}$ here is the dual version of that in [13, p. 584], which are conjugate to each other by the Atkin-Lehner involution $W_{\mathfrak{n}}$; cf. [14, p. 398].
@@
For a prime ideal $\mathfrak{q}\lhd A$ not dividing $\mathfrak{n}$, the Hecke operator $T_{\mathfrak{q}}$ commutes with the Atkin–Lehner involution $W_{\mathfrak{n}}$, so $T_{\mathfrak{q}}$ coincides with its dual; cf. [13, p. 584] and [14, p. 398].
@@
Then for a cusp $\Gamma_{0}(\mathfrak{n})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i}\end{bmatrix}$ of $X_{0}(\mathfrak{n})_{\mathbb{C}_{\infty}}$ in the form defined in Section 2, we have
$\alpha(\Gamma_{0}(\mathfrak{n})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i}\end{bmatrix})=\begin{cases}\Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i}\end{bmatrix},&\text{if $0\leq i\leq r-2$.}\\ \Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot[\infty],&\text{if $i=r-1$ or $r$.}\end{cases}$
$\beta(\Gamma_{0}(\mathfrak{n})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i}\end{bmatrix})=\begin{cases}\Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot[0],&\text{if $i=0$ or $1$.}\\ \Gamma_{0}(\mathfrak{n}/\mathfrak{p})\cdot\begin{bmatrix}a\\ \mathfrak{p}^{i-1}\end{bmatrix},&\text{if $2\leq i\leq r$.}\end{cases}$
Moreover, the degeneracy coverings $\alpha$ and $\beta$ have degree $|\mathfrak{p}|$. From the above and [9, Lemma 2.2], $\alpha$ and $\beta$ also induce maps between the divisor groups:
@@
Now, we consider $U_{\mathfrak{p}}:=\alpha^{\ast}\circ\beta_{\ast}:\operatorname{Div}(X_{0}(\mathfrak{n}))\rightarrow\operatorname{Div}(X_{0}(\mathfrak{n}))$. Recall that $(P(\mathfrak{n})_{d})$ is the sum of all the cusps of $X_{0}(\mathfrak{n})$ of height $d$. We have:
@@
In Example 2, we have the decomposition:
@@
The first statement of the lemma follows directly from [9, Lemma 2.1]. By Lemma 4.4, we have $\beta_{\ast}(\widetilde{r}(\Delta_{\mathfrak{p}})):=\widetilde{r}(\Delta_{\mathfrak{p}})|U_{\mathfrak{p}}=|\mathfrak{p}|\cdot\widetilde{r}(\Delta)$. Hence, we derive that $\beta_{\ast}$ maps $\operatorname{div}(\Delta_{\mathfrak{p}})\in\operatorname{Div}(X_{0}(\mathfrak{n}))$ to $|\mathfrak{p}|\cdot\operatorname{div}(\Delta)\in\operatorname{Div}(X_{0}(\mathfrak{n}/\mathfrak{p}))$. From [9, Section 2.1], the divisor of $\Delta_{\mathfrak{p}}$ on $X_{0}(\mathfrak{n})$ is
@@
where $b_{j}\in\mathbb{N}$. Now, the result for $2\leq j\leq r$ follows from $\beta_{\ast}(P(\mathfrak{n})_{\mathfrak{p}^{j}})=\frac{b_{j-1}}{a_{j}}\cdot(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{j-1}})$. For $j=0$ or $1$, we have $\beta_{\ast}(P(\mathfrak{n})_{1})=x\cdot(P(\mathfrak{n}/\mathfrak{p})_{1})$ and $\beta_{\ast}(P(\mathfrak{n})_{\mathfrak{p}})=y\cdot(P(\mathfrak{n}/\mathfrak{p})_{1})$ for some $x,y\in\mathbb{N}$. Consider $\widetilde{r}(\phi)\in\mathcal{H}(\mathfrak{n},\mathbb{Z})$ with
$\phi:=\left(\frac{\Delta}{\Delta_{\mathfrak{p}}}\right)^{|\mathfrak{p}|}\left(\frac{\Delta_{\mathfrak{p}^{2}}}{\Delta_{\mathfrak{p}}}\right)\in\mathcal{O}(X_{0}(\mathfrak{n}))^{\ast}.$
According to Lemma 4.3 and 4.4, we have
$\beta_{\ast}(\widetilde{r}(\phi)):=\widetilde{r}(\phi)|U_{\mathfrak{p}}=0.$
Moreover, referring to [9, Section 2.1], the divisor of $\phi$ on $X_{0}(\mathfrak{n})$ is given by:
@@
Since $\beta_{\ast}(\operatorname{div}(\phi))=0$, we have $y=\frac{|\mathfrak{p}|-1}{q-1}\cdot x$. Considering $a_{0}$, $a_{1}$, and $b_{0}$, we find
@@
Lemma 5.3 (cf. [18, Lemma 3.4]).
Fix a prime power ideal $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ with $r\geq 2$. Let $\beta_{\ast}:\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{p}^{r}))(K)\rightarrow\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{p}^{r-1}))(K)$ be the map induced from Lemma 5.2. Then
@@
Let $C_{i}:=(P(\mathfrak{n})_{\mathfrak{p}^{i}})-\deg(P(\mathfrak{n})_{\mathfrak{p}^{i}})\cdot[\infty]\in\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$ for $10\leq i\leq r-$ 1, and $C_{i}^{\prime}:=(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{i}})-\deg(P(\mathfrak{n}/\mathfrak{p})_{\mathfrak{p}^{i}})\cdot[\infty]\in\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}/\mathfrak{p}))(K)$. Then $\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$ is generated by $\{C_{0},C_{1},\cdots,C_{r-1}\}$. By Lemma 5.2, we have
@@
6 Proof of the Main Theorem
Fix a prime power ideal $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ with $r\geq 2$. We prove Theorem 1.7 by induction on $r$. It is known that $\mathcal{C}(\mathfrak{p})=\mathcal{T}(\mathfrak{p})$. Assume that $\mathcal{C}(\mathfrak{p}^{r-1})_{\ell}=\mathcal{T}(\mathfrak{p}^{r-1})_{\ell}$ for a prime $\ell\nmid q(q-1)$. We want to show that $\mathcal{C}(\mathfrak{p}^{r})_{\ell}=\mathcal{T}(\mathfrak{p}^{r})_{\ell}$. Consider the following diagram:
@@ 6 Proof of the Main Theorem
We call $U_{\mathfrak{p}}(\mathcal{T}(\mathfrak{p}^{r}))_{\ell}\subset\mathcal{T}(\mathfrak{p}^{r})_{\ell}$ the old part of $\mathcal{T}(\mathfrak{p}^{r})_{\ell}$. By Lemma 5.3 and induction hypothesis, we have
@@
From the above and [9, Lemma 3.6], the order of $\overline{C^{\prime}}$ in $\mathcal{C}(\mathfrak{p}^{r})$ divides $M(\mathfrak{p})|\mathfrak{p}|^{r-2}$. Since $\frac{\Delta_{\mathfrak{p}^{2}}}{\Delta_{\mathfrak{p}}}$ does not have a $p$-th root in $\mathcal{O}(\Omega)^{\ast}$ by [5, Corollary 3.18], we see that $|\mathfrak{p}|^{r-2}$ divides $\operatorname{ord}(\overline{C^{\prime}})$; cf. [9, Lemma 3.3]. Let $e:=\begin{pmatrix}\pi_{\infty}^{2}&\pi_{\infty}\\ 0&1\end{pmatrix}\in E(\mathscr{T})$; cf. [9, Section 2.2]. By [9, Lemma 2.4], we compute that
@@
which implies that $M(\mathfrak{p})$ also divides $\operatorname{ord}(\overline{C^{\prime}})$; cf. [9, Lemma 3.3]. Then we conclude that $\operatorname{ord}(\overline{C^{\prime}})=M(\mathfrak{p})|\mathfrak{p}|^{r-2}$.
@@
Combining the above with Theorem 1.4 and Lemma 5.3, the prime-to-$p$ part of $\mathcal{C}(\mathfrak{p}^{r})[U_{\mathfrak{p}}]$ is cyclic and generated by $|\mathfrak{p}|^{r-2}\cdot\overline{C^{\prime}}$ of order $M(\mathfrak{p})$. In contrast, the $p$-part of $\mathcal{C}(\mathfrak{p}^{r})[U_{\mathfrak{p}}]$ may not be cyclic.
@@
for a prime $\ell\nmid q(q-1)$ and large enough $k$. We claim that
$\mathcal{E}_{00}(\mathfrak{n},\mathbb{Z}/\ell^{k}\mathbb{Z})[U_{\mathfrak{p}}]\hookrightarrow\frac{\mathbb{Z}_{\ell}}{M(\mathfrak{p})\mathbb{Z}_{\ell}}.$
@@
Recall that the order of $\overline{C^{\prime}}$ in $\mathcal{C}(\mathfrak{p}^{r})$ is $M(\mathfrak{p})|\mathfrak{p}|^{r-2}$. Then by Theorem 3.4 and Equation (5), we have $E_{\mathfrak{n}}=M(\mathfrak{p})|\mathfrak{p}|^{r-2}\cdot(\widetilde{r}\circ g(C^{\prime}))\in\mathcal{H}(\mathfrak{n},\mathbb{Z})$.
@@
By Lemma 4.3 and 4.4, for any prime $\mathfrak{q}\lhd A$ not equal to $\mathfrak{p}$, we have $E_{\mathfrak{n}}|T_{\mathfrak{q}}=(|\mathfrak{q}|+1)E_{\mathfrak{n}}$ and $E_{\mathfrak{n}}|U_{\mathfrak{p}}=0$. Hence, $E_{\mathfrak{n}}$ is annihilated by both the actions of the Eisenstein ideal $\mathfrak{E}(\mathfrak{n})$ and the Hecke operator $U_{\mathfrak{p}}$. ∎
@@
This follows from Proposition 4.2 and the fact that $(f|T_{\mathfrak{m}})^{\ast}(1)=|\mathfrak{m}|f^{\ast}(\mathfrak{m})$; cf. [12, Lemma 3.2]. ∎
We provide an Atkin-Lehner type result similar to [15, Theorem 2.18]:
@@
By [15, Lemma 2.13], $h$ is $\Gamma_{0}(\mathfrak{n}/\mathfrak{p},\mathfrak{p})$-invariant and harmonic. Moreover, by Theorem 3.3, for any $e=\begin{pmatrix}\pi_{\infty}^{k}&y\\ 0&1\end{pmatrix}\in E(\mathscr{T})$ and $a\in A$, we have
@@
Note that $R\cdot E_{\mathfrak{n}}\subseteq\mathcal{E}(\mathfrak{n},R)[U_{\mathfrak{p}}]$. It suffices to prove the reverse inclusion. By Equation (2) and Lemma 3.6, we have
@@
First, by Lemma 6.4 and Theorem 6.5, since $\tilde{f}^{\ast}(1)=0$, there exists $h\in\mathcal{H}(\mathfrak{n}/\mathfrak{p},R)$ which is $\Gamma_{\infty}$-invariant such that $h|B_{\mathfrak{p}}=\tilde{f}$. Note that
@@
Second, by [13, Lemma 2.23], since $h$ is $\Gamma_{\infty}$-invariant, we have
@@
Recall that
$\left\{\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix}\in E(\mathscr{T})\mid j\geq 0\right\}$
forms an end on the Bruhat-Tits tree $\mathscr{T}$ towards $\infty$; cf. [5, pp. 184 - 185]. Then consider the action of $\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\in G(A)$ on $\mathscr{T}$. We obtain that
$\left\{e_{j}:=\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\begin{pmatrix}\pi_{\infty}^{-j}&0\\ 0&1\end{pmatrix}\in E(\mathscr{T})\mid j\geq 0\right\}$
forms another end on $\mathscr{T}$ corresponding to $\begin{bmatrix}1\\ \mathfrak{p}\end{bmatrix}=\begin{pmatrix}1&0\\ \mathfrak{p}&1\end{pmatrix}\cdot\begin{bmatrix}1\\ 0\end{bmatrix}\in\mathbb{P}^{1}(K_{\infty})$. This implies that $\{\Gamma_{0}(\mathfrak{n})\cdot e_{j}\mid j\geq 0\}$ forms an end on the quotient graph $\Gamma_{0}(\mathfrak{n})\backslash\mathscr{T}$ corresponding to the cusp $\Gamma_{0}(\mathfrak{n})\cdot\begin{bmatrix}1\\ \mathfrak{p}\end{bmatrix}$ on $X_{0}(\mathfrak{n})$; cf. [15, Definition 2.5]. Before computing $\mathcal{E}_{0}(\mathfrak{n},R)[U_{\mathfrak{p}}]$, we consider the following:
Lemma 6.7 (cf. [9, Lemma 2.4]).
@@
    $\widetilde{r}(\Delta)(e_{j})=-(q-1)(q^{j+1}-q-1)$.
@@
    $\widetilde{r}(\Delta_{\mathfrak{p}^{k}})(e_{j})=-(q-1)(q^{j+1}|\mathfrak{p}|^{2-k}-q-1).$
@@
$\widetilde{r}(\Delta)(e_{j})=\widetilde{r}(\Delta)(\begin{pmatrix}\pi_{\infty}^{-j}&0\\ \mathfrak{p}\pi_{\infty}^{-j}&1\end{pmatrix})$
@@
where the penultimate equality comes from [5, Equation (2.3)].